# uGMM-NN: Univariate Gaussian Mixture Model Neural Network

### v1

@@ 5 Discussion and Limitations
The uGMM-NN integrates probabilistic modeling into deep learning, offering interpretable uncertainty quantification through its mixture components. Compared to standard MLP neurons, each uGMM neuron introduces additional parameters, specifically, means, variances, and mixing coefficients for each input, enabling richer, multimodal representations at the cost of increased model size.
Tying the means $(\mu_{j,k})$ of each Gaussian component to the corresponding input activation $(x_{k})$, i.e., setting $\mu_{j,k}=x_{k}$, is a straightforward way to reduce the number of learnable parameters. In this case, the neuron only needs to learn the mixing coefficients $(\pi_{j,k})$ and variances $(\sigma_{j,k}^{2})$, simplifying the model while reducing expressivity, since the neuron can no longer shift submodes independently of the inputs.
@@ 6 Conclusion and Future Work
This paper introduced the Univariate Gaussian Mixture Model Neural Network (uGMM-NN), a novel feedforward architecture that integrates probabilistic reasoning directly into the computational units of neural networks. By embedding a univariate Gaussian mixture model (uGMM) into each neuron, uGMM-NN provides a richer parameterization of activations, allowing the network to capture multimodality and quantify uncertainty in a way that standard feedforward layers cannot. This makes uGMM-NN an attractive building block for architectures that require both predictive accuracy and probabilistic interpretability.
Our preliminary experiments on the Iris and MNIST benchmarks demonstrate that uGMM-NN achieves competitive performance with standard feedforward neural networks (FFNNs) when trained discriminatively. These results highlight the feasibility of replacing conventional neurons with mixture-based alternatives, offering a new design space for deep learning architectures.

### v2

@@ 5 Discussion and Limitations
The uGMM-NN integrates probabilistic modeling into deep learning, offering interpretable uncertainty quantification through its mixture components. Compared to standard MLP neurons, each uGMM neuron introduces additional parameters, specifically, means, variances, and mixing coefficients for each input, enabling richer, multimodal representations, albeit at the cost of increased model size.
Tying the means $(\mu_{j,k})$ of each Gaussian component to the corresponding input activation $(x_{k})$, i.e., setting $\mu_{j,k}=x_{k}$, is a straightforward way to reduce the number of learnable parameters. In this case, the neuron only needs to learn the mixing coefficients $(w_{j,k})$ and variances $(\sigma_{j,k}^{2})$, simplifying the model while reducing expressivity, since the neuron can no longer shift submodes independently of the inputs.
@@ 6 Conclusion and Future Work
This paper introduced the Univariate Gaussian Mixture Model Neural Network (uGMM-NN), a novel feedforward architecture that integrates probabilistic reasoning directly into the computational units of neural networks. By embedding a univariate Gaussian mixture model (uGMM) into each neuron, uGMM-NN provides a richer parameterization of activations, enabling the network to capture multimodality and quantify uncertainty in ways that standard feedforward layers cannot. This combination of predictive capability and probabilistic interpretability makes uGMM-NN a promising building block for complex neural architectures.
Our preliminary experiments demonstrate that uGMM-NN achieves competitive performance compared to standard feedforward neural networks (FFNNs) when trained discriminatively. These results highlight the feasibility of replacing conventional neurons with mixture-based alternatives, opening a new design space for deep learning architectures. Importantly, the probabilistic structure of uGMM neurons also exposes introspective signals, such as uncertainty and mixture spread, which could enable higher-level reflection and adaptive learning mechanisms.
@@
  • Investigating sparse uGMM layers, in which each neuron maintains a controlled subset of connections rather than using the full $N$-component expansion of the current design, or expanding across additional dimensions to add additional mixture components. In particular, combining sparse uGMM layers with architectures such as Transformers could enable attention-guided probabilistic representations, selectively propagating distributional information along the most relevant connections while reducing computational overhead.